given this:
F(x)= 2x - 4
G(x)= 5 + x^2
K(x) = 7
Find F(1) + G(-2) + K(0)

Answer by fractalier(6550)

You have to plug values in to their respective functions and calculate...so that for
F(x)= 2x - 4
G(x)= 5 + x^2
K(x) = 7
we have
F(1) + G(-2) + K(0) =
[2(1) - 4] + [5 + (-2)^2] + 7 =
-2 + 9 + 7 =
14